# Positivity is undecidable in tensor products of free algebras

### 09 Dec 2023 v1

@@ 4.2. Construction of $\mathcal{A}_{\mathcal{S}}(m)$
Let $\pi:\mathbb{Q}^{*}\langle x_{1},\ldots,x_{N}\rangle\to\mathbb{Q}\mathbb{Z}_{2}^{*N}$ be the homomorphism sending $x_{i}$ to the $i$th generator of $\mathbb{Z}_{2}^{*N}$. Also let $q=q_{2}\circ q_{1}$, where $q_{1}:\mathbb{Q}\mathbb{Z}_{2}^{*N}\to\mathbb{Q}\mathbb{Z}_{2}^{*\mathcal{X}}$ is the homomorphism sending $x_{i}\mapsto x_{i}$ for $1\leq i\leq n_{0}$ and $x_{i}\mapsto 1$ for $i>n_{0}$, and $q_{2}:\mathbb{Q}\mathbb{Z}_{2}^{*\mathcal{X}}\to\mathcal{A}_{\mathcal{S}}(m)$ is the quotient map. By Theorem 4.1, if $m\in\mathcal{S}$ then there is a tracial state $\tau$ on $\mathcal{A}_{\mathcal{S}}(m)$ such that $\tau(\widetilde{P}_{0}^{2})>0$. By Lemma 3.11, $\pi(t)=0$ for all $t\in\big{(}\bigcup_{r\in\mathcal{R}_{m}\cup\mathcal{R}_{m}^{*}}\mathcal{W}(r)\big{)}\cup\mathcal{W}(\widetilde{P}_{0})$. Since $q(r)=0$ for all $r\in\mathcal{R}_{m}$, $q(\pi(r^{*}r))=0$ for all $r\in\mathcal{W}_{m}\cup\mathcal{W}_{m}^{*}$, and hence $(\tau\circ q)(\pi(\alpha(m)))=-\frac{1}{\widetilde{\Lambda}^{2}m^{2k^{\prime}}}\tau(\widetilde{P}_{0}^{2})<0$. We conclude that $\pi(\alpha(m))$ is not trace-positive.
Now suppose $m\notin\mathcal{S}$, and that $\tau$ is a tracial state on *$\mathbb{C}^{*}\langle x_{1},\ldots,x_{N}\rangle$. Let $\epsilon:=\sum_{r\in\mathcal{W}_{m}\cup\mathcal{W}_{m}^{*}}\tau(r^{*}r)$. Then $\tau$ is a tracial $(\epsilon,\mathcal{W}_{m})$-state, so by Proposition 5.4, $\tau(\widetilde{P}^{*}_{0}\widetilde{P}_{0})=\lVert\widetilde{P}_{0}\rVert^{2}_{\tau}\leq\widetilde{\Lambda}^{2}m^{2k^{\prime}}\epsilon$. Hence $\tau(\alpha(m))\geq 0$. Thus the complement $\overline{\mathcal{S}}$ of $\mathcal{S}$ reduces to the promise problem in the theorem statement. Taking $\mathcal{S}$ to be RE-hard, we see that the promise problem is coRE-hard. ∎
@@
Suppose either $N\geq 2$ and $\mathcal{C}=\mathbb{C}\mathbb{Z}_{3}^{*N}$, or $N\geq 3$ and $\mathcal{C}=\mathbb{C}\mathbb{Z}_{2}^{*N}$. Let $\mathcal{P}=\mathbb{Q}^{*}\langle x_{1},\ldots,x_{N}\rangle$, and let $\pi:\mathcal{P}\to\mathcal{C}$ be the homomorphism sending $x_{i}$ to the $i$th generator of $\mathcal{C}$. Then the promise problem
@@
To get Theorem 1.1 from Theorem 6.1, suppose that $N$, $\mathcal{C}$, $\mathcal{P}$, and $\pi$ are as in the theorem, and that $\pi\otimes\pi$ factors through a ***-algebra $\mathcal{A}$, meaning that there are homomorphisms $\pi_{1}:\mathcal{P}\otimes\mathcal{P}\to\mathcal{A}$ and $\pi_{2}:\mathcal{A}\to\mathcal{C}\otimes\mathcal{C}$ such that $\pi\otimes\pi=\pi_{2}\circ\pi_{1}$. Suppose also that the first homomorphism $\pi_{1}$ in this factorization is computable. Then:
@@
The proof is the same as the proof of 5.2, with the word “trace” deleted. Theorem 1.1 follows immediately from 6.2 by taking $\mathcal{A}$ to be one of *$\mathbb{Q}^{*}\langle x_{1},\ldots,x_{N}\rangle\otimes\mathbb{Q}^{*}\langle x_{1},\ldots,x_{N}\rangle$ or $\mathbb{Q}\mathcal{F}_{N}\otimes\mathbb{Q}\mathcal{F}_{N}$ with $N\geq 2$, or $\mathbb{Q}\mathbb{Z}_{M}^{*N}\otimes\mathbb{Q}\mathbb{Z}_{M}^{*N}$ with $N\geq 3,M\geq 2$ or $N\geq 2,M\geq 3$. As with 5.2, 6.2 also applies to other algebras such as the algebra of contractions and the algebra of self-adjoint contractions, as well as free products of any of the algebras mentioned.
For the proof of Theorem 6.1, we use the same setup as in the previous section: we suppose $\mathcal{S}\in\mathbb{N}$ is an RE set, and let $\mathcal{X},\widetilde{P}_{n},\widetilde{X}_{n}$, $n\geq 0$, and $\mathcal{A}_{\mathcal{S}}(m)=\mathbb{C}^{*}\langle\mathbb{C}\mathbb{Z}_{2}^{*\mathcal{X}}:\mathcal{R}_{m}\rangle$, $m\in\mathbb{N}$, be as in Theorem 4.1. In addition, let $\iota:\mathbb{C}\mathbb{Z}_{2}^{*\mathcal{X}}\rightarrow\mathbb{C}\mathbb{Z}_{2}^{*\mathcal{X}}\otimes\mathbb{C}\mathbb{Z}_{2}^{*\mathcal{X}}:a\mapsto a\otimes 1$ be the left inclusion.
@@
There are positive integers $\Gamma,k$ such that if $m\notin\mathcal{S}$, $\varphi$ is an $(\epsilon,\mathcal{X})$-synchronous state on $\mathbb{C}\mathbb{Z}_{2}^{*\mathcal{X}}\otimes\mathbb{C}\mathbb{Z}_{2}^{*\mathcal{X}}$, and $\tau:=\varphi\circ\iota$ is an $(\epsilon,\mathcal{R}_{m})$-state on $\mathbb{C}\mathbb{Z}_{2}^{*\mathcal{X}}$ for some $\epsilon\geq 0$, then $\lVert\widetilde{P}_{0}\rVert_{\tau}\leq\Gamma m^{k}\sqrt{\epsilon}$.
@@
Suppose $m\notin\mathcal{S}$, $\epsilon\geq 0$, and $\varphi$ is an $(\epsilon,\mathcal{X})$-synchronous state on $\mathbb{C}\mathbb{Z}_{2}^{*\mathcal{X}}\otimes\mathbb{C}\mathbb{Z}_{2}^{*\mathcal{X}}$ such that $\tau:=\varphi\circ\iota$ is an $(\epsilon,\mathcal{R}_{m})$-state on $\mathbb{C}\mathbb{Z}_{2}^{*\mathcal{X}}$. By Theorem 4.1, part (a), $\widetilde{P}_{n}$ is a contraction in $\mathbb{C}\mathbb{Z}_{2}^{*\mathcal{X}}$, $\widetilde{X}_{n}$ is a monomial in *$\mathbb{C}^{*}\langle\mathcal{X}\rangle$ of degree $14n+$ 1, and $\widetilde{X}_{n}$ is a unitary involution in $\mathbb{C}\mathbb{Z}_{2}^{*\mathcal{X}}$, so Lemma 3.7 implies
@@
There are positive integers $\widetilde{\Gamma},k^{\prime}$ such that if $m\notin\mathcal{S}$, $\varphi$ is an $(\epsilon,\widetilde{\mathcal{W}}_{m})$-state on *$\mathbb{C}^{*}\langle\mathcal{X}\rangle\otimes\mathbb{C}^{*}\langle\mathcal{X}\rangle$, and $\varphi$ is $(\epsilon,\mathcal{X})$-synchronous, then $\lVert\widetilde{P}_{0}\otimes 1\rVert_{\varphi}\leq\widetilde{\Gamma}m^{k^{\prime}}\sqrt{\epsilon}$.
@@
Let $\Gamma,k$ be integers from Proposition 6.3, and let $D$ be the integer from Theorem 4.1, part (d). Suppose $m\not\in\mathcal{S}$, $\epsilon\geq 0$, and that $\varphi$ is an $(\epsilon,\widetilde{\mathcal{W}}_{m})$-state on *$\mathbb{C}^{*}\langle\mathcal{X}\rangle\otimes\mathbb{C}^{*}\langle\mathcal{X}\rangle$ which is $(\epsilon,\mathcal{X})$-synchronous. Since $\varphi(s^{*}s\otimes 1)\leq\epsilon$ and $\varphi(1\otimes s^{*}s)\leq\epsilon$ for all $s\in\bigcup_{x\in\mathcal{X}}\mathcal{W}(x)$, and $\{s\otimes 1:s\in\bigcup_{r\in\mathcal{R}_{m}\cup\mathcal{R}_{m}^{*}}\mathcal{W}(r)\}\subseteq\widetilde{\mathcal{W}}_{m}$, by Lemma 3.12, there is a $(25\epsilon,\mathcal{X})$-synchronous state $\widehat{\varphi}$ on $\mathbb{C}\mathbb{Z}_{2}^{*\mathcal{X}}\otimes\mathbb{C}\mathbb{Z}_{2}^{*\mathcal{X}}$ such that
@@
for all $r\in\mathcal{R}_{m}\cup\mathcal{R}_{m}^{*}$. Since $\{r\otimes 1:r\in\mathcal{R}_{m}\}\subseteq\widetilde{\mathcal{W}}_{m}$, for every $r\in\mathcal{R}_{m}\cup\mathcal{R}_{m}^{*}$, $\lVert r\otimes 1\rVert_{\varphi}\leq\sqrt{\epsilon}$, and hence $\lVert r\otimes 1\rVert_{\widehat{\varphi}}\leq(2Dm+1)\sqrt{\epsilon}$. Rounding $12Dm+$ 1 up to $32Dm+$ 3 so that $(2Dm+3)^{2}\geq 25$, we conclude that $\widehat{\varphi}$ is a $((2Dm+3)^{2}\epsilon,\mathcal{X})$-synchronous state on $\mathbb{C}\mathbb{Z}_{2}^{*\mathcal{X}}\otimes\mathbb{C}\mathbb{Z}_{2}^{*\mathcal{X}}$ and $\widehat{\varphi}\circ\iota$ is a $((2Dm+3)^{2}\epsilon,\mathcal{R}_{m})$-state on $\mathbb{C}\mathbb{Z}_{2}^{*\mathcal{X}}$. By Proposition 6.3, $\lVert\widetilde{P}_{0}\otimes 1\rVert_{\widehat{\varphi}}\leq\Gamma m^{k}\sqrt{(2Dm+3)^{2}\epsilon}=(2Dm+3)\Gamma m^{k}\sqrt{\epsilon}$. Again by Lemma 3.12, since $\{s\otimes 1:s\in\mathcal{W}(\widetilde{P}_{0})\}\subseteq\widetilde{\mathcal{W}}_{m}$, we have
@@
Let $(\pi,v)$ be the GNS representation of $\tau$, and let $\mathcal{C}=\overline{\pi(\mathcal{A})}$ be the closure of $\pi(\mathcal{A})$ in the operator norm, where $\mathcal{A}=\mathbb{C}\mathbb{Z}_{2}^{*\mathcal{X}}$. By [BO08, Proposition 6.1.2], the tracial state defined by $v$ on $\mathcal{C}$ extends to a state $\varphi^{\prime}:\mathcal{C}\otimes\mathcal{C}^{op}\to\mathbb{C}:a\otimes b\mapsto\langle v,abv\rangle$. Note that the isomorphism $\omega_{\mathcal{X}}:\mathbb{C}^{*}\langle\mathcal{X}\rangle\to\mathbb{C}^{*}\langle\mathcal{X}\rangle^{op}$ defined in Section 3 descends to an isomorphism $\mathcal{A}\to\mathcal{A}^{op}$, and that $\pi$ defines a ***-homomorphism $\pi^{op}:\mathcal{A}^{op}\to\mathcal{C}^{op}$. Let $\varphi=\varphi^{\prime}\circ(\pi\otimes(\pi^{op}\circ\omega_{\mathcal{X}}))$. Then $\varphi$ is a state on $\mathcal{A}\otimes\mathcal{A}$ with $\varphi(a\otimes b)=\langle v,\pi(a\omega_{\mathcal{X}}(b))v\rangle=\tau(a\cdot\omega_{\mathcal{X}}(b))$. Since
@@
The remaining two lemmas are used to get from $\mathbb{Q}\mathcal{F}_{N}$ and $\mathbb{Q}\mathbb{Z}_{2}^{*N}$ for large $N$ down to $\mathbb{Q}\mathbb{Z}_{3}^{*2}$ and $\mathbb{Q}\mathbb{Z}_{2}^{*3}$.
@@
One way to prove this lemma is to observe that $\mathbb{C}H$ and $\mathbb{C}G$ are subalgebras of the full group $C^{*}$-algebras $C^{*}(H)$ and $C^{*}(G)$. By definition of the full group $C^{*}$-algebra, any state on $\mathbb{C}H$ is the restriction of a state on $C^{*}(H)$, and since $C^{*}(H)$ is a subalgebra of $C^{*}(G)$, any state on $C^{*}(H)$ extends to a state on $C^{*}(G)$. There is also a ***-algebra proof which we expect is known to experts, but have been unable to find in the literature, so we include a proof for reference.
Proof of Lemma 6.6.
Let $\mathbb{C}H_{+}$ and $\mathbb{C}G_{+}$ be the convex cones of sums of squares in $\mathbb{C}H$ and $\mathbb{C}G$ respectively. We will show that $\mathbb{C}G_{+}\cap\mathbb{C}H=\mathbb{C}H_{+}$. Indeed, let $r:\mathbb{C}G\to\mathbb{C}H$ be the linear function with $r(g)=g$ if $g\in H$ and $r(g)=0$ if $g\not\in H$. Pick a set of representatives $X\subset G$ for the coset space $G/H$. Any element $\alpha$ of $\mathbb{C}G$ can be written as $\sum_{x\in X}x\alpha_{x}$, where $\alpha_{x}\in\mathbb{C}H$, and $\alpha_{x}$ is zero for all but finitely many elements $x\in X$. If $h,h^{\prime}\in H$ and $x,y\in X$, then $h^{-1}x^{-1}yh^{\prime}\in H$ if and only if $x=y$, so
$r(\alpha^{*}\alpha)=\sum_{x,y}r(\alpha_{x}^{*}x^{-1}y\alpha_{y})=\sum_{x}\alpha_{x}^{*}\alpha_{x}$
is a sum of squares in $\mathbb{C}H$. Consequently, if $a\in\mathbb{C}H$ and there are $\alpha_{1},\ldots,\alpha_{k}\in\mathbb{C}G$ such that $a=\sum_{i}\alpha_{i}^{*}\alpha_{i}$, then $a=r(a)=\sum_{i}r(\alpha_{i}^{*}\alpha_{i})$ is a sum of squares in $\mathbb{C}H$, so $\mathbb{C}G_{+}\cap\mathbb{C}H=\mathbb{C}H_{+}$ as desired.
If $f$ is a state on $\mathbb{C}H$, then $f$ is nonnegative on $\mathbb{C}H\cap\mathbb{C}G_{+}=\mathbb{C}H_{+}$. Since $\mathbb{C}G$ is archimedean, the M. Riesz extension theorem implies that $f$ extends to a real linear functional on $\text{span}_{\mathbb{R}}\mathbb{C}G_{+}$ which is nonnegative on $\mathbb{C}G_{+}$. From there, we can extend further to get a linear functional $\widetilde{f}$ on $\mathbb{C}G$ such that $\widetilde{f}(a^{*})=\overline{\widetilde{f}(a)}$ for all $a\in\mathbb{C}G$, $\widetilde{f}|_{\mathbb{C}H}=f$, and $\widetilde{f}$ is nonnegative on $\mathbb{C}G_{+}$. Since $\widetilde{f}(1)=f(1)=1$, the archimedean condition implies that $\widetilde{f}$ is a state. ∎
@@
Suppose $N\geq 1$ and $G$ is either $\mathcal{F}_{K}$ or $\mathbb{Z}_{3}^{*K}$ with $K\geq 2$, or $\mathbb{Z}_{2}^{*K}$ with $K\geq 3$. Then there is an injective group homomorphism $\mu:\mathcal{F}_{N}\to G$ and a computable ***-homomorphism $\widetilde{\mu}:\mathbb{Q}^{*}\langle x_{1},\ldots,x_{N}\rangle\to\mathbb{Q}^{*}\langle x_{1},\ldots,x_{K}\rangle$ such that the diagram

### 18 Mar 2026 v2

@@ 4.2. Construction of $\mathcal{A}_{\mathcal{L}}(m)$
Let $\pi:\mathbb{Q}^{*}\langle x_{1},\ldots,x_{N}\rangle\to\mathbb{Q}\mathbb{Z}_{2}^{*N}$ be the homomorphism sending $x_{i}$ to the $i$th generator of $\mathbb{Z}_{2}^{*N}$. Also let $q=q_{2}\circ q_{1}$, where $q_{1}:\mathbb{Q}\mathbb{Z}_{2}^{*N}\to\mathbb{Q}\mathbb{Z}_{2}^{*\mathcal{X}}$ is the homomorphism sending $x_{i}\mapsto x_{i}$ for $1\leq i\leq n_{0}$ and $x_{i}\mapsto 1$ for $i>n_{0}$, and $q_{2}:\mathbb{Q}\mathbb{Z}_{2}^{*\mathcal{X}}\to\mathcal{A}_{\mathcal{L}}(m)$ is the quotient map. By Theorem 4.1, if $m\in\mathcal{L}$ then there is a tracial state $\tau$ on $\mathcal{A}_{\mathcal{L}}(m)$ such that $\tau(\widetilde{P}_{0}^{2})>0$. By Lemma 3.11, $\pi(t)=0$ for all $t\in\big(\bigcup_{r\in\mathcal{R}_{m}\cup\mathcal{R}_{m}^{*}}\mathcal{W}(r)\big)\cup\mathcal{W}(\widetilde{P}_{0})$. Since $q(r)=0$ for all $r\in\mathcal{R}_{m}$, $q(\pi(r^{*}r))=0$ for all $r\in\mathcal{W}_{m}\cup\mathcal{W}_{m}^{*}$, and hence $(\tau\circ q)(\pi(\alpha(m)))=-\frac{1}{\widetilde{\Lambda}^{2}m^{2k^{\prime}}}\tau(\widetilde{P}_{0}^{2})<0$. We conclude that $\pi(\alpha(m))$ is not trace-positive.
Now suppose $m\notin\mathcal{L}$, and that $\tau$ is a tracial state on $\mathbb{C}^{*}\langle x_{1},\ldots,x_{N}\rangle$. Let $\epsilon:=\sum_{r\in\mathcal{W}_{m}\cup\mathcal{W}_{m}^{*}}\tau(r^{*}r)$. Then $\tau$ is a tracial $(\epsilon,\mathcal{W}_{m})$-state, so by Proposition 5.4, $\tau(\widetilde{P}^{*}_{0}\widetilde{P}_{0})=\lVert\widetilde{P}_{0}\rVert^{2}_{\tau}\leq\widetilde{\Lambda}^{2}m^{2k^{\prime}}\epsilon$. Hence $\tau(\alpha(m))\geq 0$. Thus the complement $\overline{\mathcal{L}}$ of $\mathcal{L}$ reduces to the promise problem in the theorem statement. Taking $\mathcal{L}$ to be RE-hard, we see that the promise problem is coRE-hard. ∎
@@
Suppose either $N\geq 2$ and $\mathcal{C}=\mathbb{Q}\mathbb{Z}_{3}^{*N}$, or $N\geq 3$ and $\mathcal{C}=\mathbb{Q}\mathbb{Z}_{2}^{*N}$. Let $\mathcal{P}=\mathbb{Q}^{*}\langle x_{1},\ldots,x_{N}\rangle$, and let $\pi:\mathcal{P}\to\mathcal{C}$ be the homomorphism sending $x_{i}$ to the $i$th generator of $\mathcal{C}$. Then the promise problem
@@
To get Theorem 1.1 from Theorem 6.1, suppose that $N$, $\mathcal{C}$, $\mathcal{P}$, and $\pi$ are as in the theorem, and that $\pi\otimes\pi$ factors through a *-algebra $\mathcal{A}$, meaning that there are homomorphisms $\pi_{1}:\mathcal{P}\otimes\mathcal{P}\to\mathcal{A}$ and $\pi_{2}:\mathcal{A}\to\mathcal{C}\otimes\mathcal{C}$ such that $\pi\otimes\pi=\pi_{2}\circ\pi_{1}$. Suppose also that the first homomorphism $\pi_{1}$ in this factorization is computable. Then:
@@
The proof is the same as the proof of 5.2, with the word “trace” deleted. Theorem 1.1 follows immediately from 6.2 by taking $\mathcal{A}$ to be one of $\mathbb{Q}^{*}\langle x_{1},\ldots,x_{N}\rangle\otimes\mathbb{Q}^{*}\langle x_{1},\ldots,x_{N}\rangle$ or $\mathbb{Q}\mathcal{F}_{N}\otimes\mathbb{Q}\mathcal{F}_{N}$ with $N\geq 2$, or $\mathbb{Q}\mathbb{Z}_{M}^{*N}\otimes\mathbb{Q}\mathbb{Z}_{M}^{*N}$ with $N\geq 3,M\geq 2$ or $N\geq 2,M\geq 3$. As with 5.2, 6.2 also applies to other algebras such as the algebra of contractions and the algebra of self-adjoint contractions, as well as free products of any of the algebras mentioned.
For the proof of Theorem 6.1, we use the same setup as in the previous section: we suppose $\mathcal{L}\subseteq\mathbb{N}$ is an RE set, and let $\mathcal{X},\widetilde{P}_{n},\widetilde{X}_{n}$, $n\geq 0$, and $\mathcal{A}_{\mathcal{L}}(m)=\mathbb{C}^{*}\langle\mathbb{C}\mathbb{Z}_{2}^{*\mathcal{X}}:\mathcal{R}_{m}\rangle$, $m\in\mathbb{N}$, be as in Theorem 4.1. In addition, let $\iota:\mathbb{C}\mathbb{Z}_{2}^{*\mathcal{X}}\rightarrow\mathbb{C}\mathbb{Z}_{2}^{*\mathcal{X}}\otimes\mathbb{C}\mathbb{Z}_{2}^{*\mathcal{X}}:a\mapsto a\otimes 1$ be the left inclusion.
@@
There are positive integers $\Gamma,k$ such that if $m\notin\mathcal{L}$, $\varphi$ is an $(\epsilon,\mathcal{X})$-synchronous state on $\mathbb{C}\mathbb{Z}_{2}^{*\mathcal{X}}\otimes\mathbb{C}\mathbb{Z}_{2}^{*\mathcal{X}}$, and $\tau:=\varphi\circ\iota$ is an $(\epsilon,\mathcal{R}_{m})$-state on $\mathbb{C}\mathbb{Z}_{2}^{*\mathcal{X}}$ for some $\epsilon\geq 0$, then $\lVert\widetilde{P}_{0}\rVert_{\tau}\leq\Gamma m^{k}\sqrt{\epsilon}$.
@@
Suppose $m\notin\mathcal{L}$, $\epsilon\geq 0$, and $\varphi$ is an $(\epsilon,\mathcal{X})$-synchronous state on $\mathbb{C}\mathbb{Z}_{2}^{*\mathcal{X}}\otimes\mathbb{C}\mathbb{Z}_{2}^{*\mathcal{X}}$ such that $\tau:=\varphi\circ\iota$ is an $(\epsilon,\mathcal{R}_{m})$-state on $\mathbb{C}\mathbb{Z}_{2}^{*\mathcal{X}}$. By Theorem 4.1, part (a), $\widetilde{P}_{n}$ is a contraction in $\mathbb{C}\mathbb{Z}_{2}^{*\mathcal{X}}$, $\widetilde{X}_{n}$ is a monomial in $\mathbb{C}^{*}\langle\mathcal{X}\rangle$ of degree $4n+1$, and $\widetilde{X}_{n}$ is a unitary involution in $\mathbb{C}\mathbb{Z}_{2}^{*\mathcal{X}}$, so Lemma 3.7 implies
@@
There are positive integers $\widetilde{\Gamma},k^{\prime}$ such that if $m\notin\mathcal{L}$, $\varphi$ is an $(\epsilon,\widetilde{\mathcal{W}}_{m})$-state on $\mathbb{C}^{*}\langle\mathcal{X}\rangle\otimes\mathbb{C}^{*}\langle\mathcal{X}\rangle$, and $\varphi$ is $(\epsilon,\mathcal{X})$-synchronous, then $\lVert\widetilde{P}_{0}\otimes 1\rVert_{\varphi}\leq\widetilde{\Gamma}m^{k^{\prime}}\sqrt{\epsilon}$.
@@
Let $\Gamma,k$ be integers from Proposition 6.3, and let $D$ be the integer from Theorem 4.1, part (d). Suppose $m\not\in\mathcal{L}$, $\epsilon\geq 0$, and that $\varphi$ is an $(\epsilon,\widetilde{\mathcal{W}}_{m})$-state on $\mathbb{C}^{*}\langle\mathcal{X}\rangle\otimes\mathbb{C}^{*}\langle\mathcal{X}\rangle$ which is $(\epsilon,\mathcal{X})$-synchronous. Since $\varphi(s^{*}s\otimes 1)\leq\epsilon$ and $\varphi(1\otimes s^{*}s)\leq\epsilon$ for all $s\in\bigcup_{x\in\mathcal{X}}\mathcal{W}(x)$, and $\{s\otimes 1:s\in\bigcup_{r\in\mathcal{R}_{m}\cup\mathcal{R}_{m}^{*}}\mathcal{W}(r)\}\subseteq\widetilde{\mathcal{W}}_{m}$, by Lemma 3.12, there is a $(25\epsilon,\mathcal{X})$-synchronous state $\widehat{\varphi}$ on $\mathbb{C}\mathbb{Z}_{2}^{*\mathcal{X}}\otimes\mathbb{C}\mathbb{Z}_{2}^{*\mathcal{X}}$ such that
@@
for all $r\in\mathcal{R}_{m}\cup\mathcal{R}_{m}^{*}$. Since $\{r\otimes 1:r\in\mathcal{R}_{m}\}\subseteq\widetilde{\mathcal{W}}_{m}$, for every $r\in\mathcal{R}_{m}\cup\mathcal{R}_{m}^{*}$, $\lVert r\otimes 1\rVert_{\varphi}\leq\sqrt{\epsilon}$, and hence $\lVert r\otimes 1\rVert_{\widehat{\varphi}}\leq(2Dm+1)\sqrt{\epsilon}$. Rounding $2Dm+1$ up to $2Dm+3$ so that $(2Dm+3)^{2}\geq 25$, we conclude that $\widehat{\varphi}$ is a $((2Dm+3)^{2}\epsilon,\mathcal{X})$-synchronous state on $\mathbb{C}\mathbb{Z}_{2}^{*\mathcal{X}}\otimes\mathbb{C}\mathbb{Z}_{2}^{*\mathcal{X}}$ and $\widehat{\varphi}\circ\iota$ is a $((2Dm+3)^{2}\epsilon,\mathcal{R}_{m})$-state on $\mathbb{C}\mathbb{Z}_{2}^{*\mathcal{X}}$. By Proposition 6.3, $\lVert\widetilde{P}_{0}\otimes 1\rVert_{\widehat{\varphi}}\leq\Gamma m^{k}\sqrt{(2Dm+3)^{2}\epsilon}=(2Dm+3)\Gamma m^{k}\sqrt{\epsilon}$. Again by Lemma 3.12, since $\{s\otimes 1:s\in\mathcal{W}(\widetilde{P}_{0})\}\subseteq\widetilde{\mathcal{W}}_{m}$, we have
@@
Let $(\pi,v)$ be the GNS representation of $\tau$, and let $\mathcal{C}=\overline{\pi(\mathcal{A})}$ be the closure of $\pi(\mathcal{A})$ in the operator norm, where $\mathcal{A}=\mathbb{C}\mathbb{Z}_{2}^{*\mathcal{X}}$. By [BO08, Proposition 6.1.2], the tracial state defined by $v$ on $\mathcal{C}$ extends to a state $\varphi^{\prime}:\mathcal{C}\otimes\mathcal{C}^{op}\to\mathbb{C}:a\otimes b\mapsto\langle v,abv\rangle$. Note that the isomorphism $\omega_{\mathcal{X}}:\mathbb{C}^{*}\langle\mathcal{X}\rangle\to\mathbb{C}^{*}\langle\mathcal{X}\rangle^{op}$ defined in Section 3 descends to an isomorphism $\mathcal{A}\to\mathcal{A}^{op}$, and that $\pi$ defines a *-homomorphism $\pi^{op}:\mathcal{A}^{op}\to\mathcal{C}^{op}$. Let $\varphi=\varphi^{\prime}\circ(\pi\otimes(\pi^{op}\circ\omega_{\mathcal{X}}))$. Then $\varphi$ is a state on $\mathcal{A}\otimes\mathcal{A}$ with $\varphi(a\otimes b)=\langle v,\pi(a\omega_{\mathcal{X}}(b))v\rangle=\tau(a\cdot\omega_{\mathcal{X}}(b))$. Since
@@
The remaining two lemmas are used to get from $\mathbb{Q}\mathcal{F}_{N}$ and $\mathbb{Q}\mathbb{Z}_{2}^{*N}$ for large $N$ down to $\mathbb{Q}\mathbb{Z}_{3}^{*2}$ and $\mathbb{Q}\mathbb{Z}_{2}^{*3}$. Both lemmas will be familiar to experts.
@@
Proof.
Let $\mathbb{C}H_{+}$ and $\mathbb{C}G_{+}$ be the convex cones of sums of squares in $\mathbb{C}H$ and $\mathbb{C}G$ respectively, and let $r:\mathbb{C}G\to\mathbb{C}H$ be the linear function with $r(g)=g$ if $g\in H$ and $r(g)=0$ if $g\not\in H$. By [SS11, Proposition 4], $r(\mathbb{C}G_{+})\subseteq\mathbb{C}H_{+}$. Consequently, if $a\in\mathbb{C}G_{+}\cap\mathbb{C}H$, then $a=r(a)\in\mathbb{C}H_{+}$, so $\mathbb{C}G_{+}\cap\mathbb{C}H=\mathbb{C}H_{+}$. If $f$ is a state on $\mathbb{C}H$, then $f$ is nonnegative on $\mathbb{C}H_{+}$. Since $\mathbb{C}G$ is archimedean, the M. Riesz extension theorem implies that $f$ extends to a real linear functional on $\text{span}_{\mathbb{R}}\mathbb{C}G_{+}$ which is nonnegative on $\mathbb{C}G_{+}$. From there, we can extend further to get a linear functional $\widetilde{f}$ on $\mathbb{C}G$ such that $\widetilde{f}(a^{*})=\overline{\widetilde{f}(a)}$ for all $a\in\mathbb{C}G$, $\widetilde{f}|_{\mathbb{C}H}=f$, and $\widetilde{f}$ is nonnegative on $\mathbb{C}G_{+}$. Since $\widetilde{f}(1)=f(1)=1$, the archimedean condition implies that $\widetilde{f}$ is a state. ∎
Another way to prove this lemma is to observe that $\mathbb{C}H$ and $\mathbb{C}G$ are subalgebras of the full group $C^{*}$-algebras $C^{*}(H)$ and $C^{*}(G)$. By definition of the full group $C^{*}$-algebra, any state on $\mathbb{C}H$ is the restriction of a state on $C^{*}(H)$, and since $C^{*}(H)$ is a subalgebra of $C^{*}(G)$, any state on $C^{*}(H)$ extends to a state on $C^{*}(G)$.
@@
Suppose $N\geq 1$ and $G$ is either $\mathcal{F}_{K}$ or $\mathbb{Z}_{3}^{*K}$ with $K\geq 2$, or $\mathbb{Z}_{2}^{*K}$ with $K\geq 3$. Then there is an injective group homomorphism $\mu:\mathcal{F}_{N}\to G$ and a computable *-homomorphism $\widetilde{\mu}:\mathbb{Q}^{*}\langle x_{1},\ldots,x_{N}\rangle\to\mathbb{Q}^{*}\langle x_{1},\ldots,x_{K}\rangle$ such that the diagram